a grandfather is 5 times as old as his granddaughter. in 6 years, he will be 4 times as old as his granddaughter will be. find their present ages.
Answer by mananth(16949)

a grandfather is 5 times as old as his granddaughter. in 6 years, he will be 4 times as old as his granddaughter will be. find their present ages.
..
grand daughter age be x
granfather age = 5x
..
after 6 years
their ages will be
granddaughter = x+6
grandfather = 5x+6
..
5x+6 = 4(x+6)
5x+6=4x+24
5x-4x=24-6
x=18 grandson's age
grandfather's age= 5x = 90 years
...
m.ananth@hotmail.ca
...